Tags:
create new tag
view all tags

2012/04/09 Calibration shift

  • Tuned all BPMs
  • Trained Laurie, Laura, Jackie, Naito and Kubo in BPM operations
  • SVD calibration (static log 50, cal step 50, +-150 non-attenuated, +-250 attenuated)
  • Check zip file logging
  • Mover control to trim only (for more stable positions)
  • Fixed S-band reference negative sign in tune script
  • Calibration for MFB2FF (now laserwire)
  • Race condition on lsis8 (qd4aff, qf3ff, qd2bff, qd2aff)
  • Fixed by added cables to clocks on lsis8 and lsis10. This messed up FB2FF, so needs recalibration (6:30am)

  • Tune IP references
  • Online resolution script in normal startup
  • Better IP BPM tuning code
  • Calibrate references to charge
  • IPBPM tune and calibration
  • Logger for DBPM diode
  • 20120410_shift.txt: 20120410_shift.txt
  • Archive of the week is bari_20120411_052719
  • Data sync to UK on 2012/04/13
  • SVN Commit 2012/04/13

Found diamond cavity!

Channel swapping went wrong, signals are in lsis10:waveform2, lsis10:waveform3 Will move to lsis13:waveform6 lsis9:waveform3

Edit | Attach | Watch | Print version | History: r8 < r7 < r6 < r5 < r4 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r8 - 13 Apr 2012 - StewartBoogert

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback